I need help

Bug Population changes when there is a breeding preference for blue rimmed bugs

Data Table - Enter your Final Bug Counts

BB Bug

Count

Bb Bug

Count

bb Bug

Count

20

0

0

                       

Percentage Tables - Enter the Final Bug percentages

  1. : Bug Type Percentage = 100% ´ (Bug Type Count) / (Total Number of Bugs)

BB Bug

Percentage

Bb Bug

Percentage

bb Bug

Percentage

100% 0 0

Tip: Blue Rimmed Baby Bug Percentage = BB Bug Percent + Bb Bug Percent

Blue Rimmed Bug

Percentage

Yellow Rimmed Bug

Percentage

100% 0

Observations and Questions

[16] Describe the bug population change results during this data run in terms of genotypes and phenotypes.

[17] Do your results suggest anything about what the composition of this population might be at some distant point in the future? Defend your answer.

[18] Based on the initial starting population, use the Hardy-Weinberg equation to predict the future bug population phenotype composition. Hint: Under the Background tab, go to the Summary of Formulas Needed for Calculations section, see the example titled Using the Hardy-Weinberg Equation, then do Step 1 and Step 2 using the initial starting population for this data run.

[19] Is this population consistent with the expectations of the Hardy-Weinberg model, that is, is this population stable? Hint: Under the Background tab, go to the Summary of Formulas Needed for Calculations section, see the example titled Using the Hardy-Weinberg Equation, then do Step 3 and Step 4 for this data run.
Answer #1

Answer 16)

The bug population has only single genotype that is BB and encodes blue phenotype. That's why there is found only single trait in the bug population. There is found blue phenotype and BB genotype.

Answer 17)

The population have 100% BB genotype and blue phenotype at some distant in the future. Because the individual in a population when undergoes breeding, then there is found only B gametes by the individuals and their fusion produces again BB individuals having blue phenotype.

Answer 18)

According to hardyweiberg equation,

Genotype frequency of B is 100/100=1

The alleleic frequency of B is calculated by

=Genotype frequency of { BB +(1/2) Bb)}

= 1 + (1/2) ×0

=1

Alleleic frequency of b is = 1- B = 1 - 1 =0

Blue phenotype percentage in future generations= BB + Bb

BB= (B) ^2 ×100= (1) ^2 ×100=100%

Bb= 2×B×b ×100 = 2×1×0×100=0%

That's why blue phenotype percentage is 100+0=100%

Yellow phenotype is = (B) ^2= (0) ^2×100=0

Answer 19)

No, this population is not consistent with the Hardy weinberg model and this population is not stable.

Because there is found very small population size. There is found genetic drift which leads to random fixation of alleles. In this case, the B gene is fixed in the bug population.